Rationalize the denominator of √27t^2r / √3tr
Assume that all variables are positive.
1
Expert's answer
2012-02-07T09:48:31-0500
sqrt(27 t^2 r)/sqrt(3 tr)=sqrt(27 t^2 r)*sqrt(3 tr)/(3tr)=
sqrt(27 t^2 r*3 tr)/(3tr) =3 sqrt(3 t^2 r*3 tr)/(3tr)=
sqrt(9 t^3 r^2)/(tr)=
=3 tr sqrt(t) /(tr)=3 sqrt(t)